OpenStudy (anonymous):

What is the common ratio between the successive height values of ball 1? Ball 2? Ball 3? ball 1: (1,84)(2,29)(3,15)(4,7)(5,2) ball 2: (1,84)(2,39)(3,19)(2,8)(3,4) ball 3: (1,84)(2,59)(3,43)(4,36)(5,27)

OpenStudy (anonymous):

I had to do an activity where I measured the height of a ball when I dropped it from 7 feet high

OpenStudy (anonymous):

please someone help me :'(

OpenStudy (anonymous):

Thinking...

OpenStudy (anonymous):

okay thankyou soo much!

OpenStudy (anonymous):

Aha. Common ratio= h2/h1 where h1=any height. h2=height of the next bounce. Note that since this is an experiment, the common ratio will differ a little depending upon which two consecutive heights you choose.

OpenStudy (anonymous):

so I just take the height divided by the next height so 84/29 = 2.9

OpenStudy (anonymous):

And I do that for each of the numbers?

OpenStudy (anonymous):

wait what do I write then? Which one? cause they're all different

OpenStudy (anonymous):

No, it's 29/84. & yea that's what I said. They will not match exactly because of experimental errors. But they should be close nevertheless.

OpenStudy (anonymous):

so 29/84=0.34 so thats my answer? or do I write all of the numbers I get?

OpenStudy (anonymous):

Any one will do.

OpenStudy (anonymous):

okay thankyou soo muchh!

OpenStudy (anonymous):

yw. :)

OpenStudy (anonymous):

can anyone answer this? If ball 1 were dropped from 2 feet higher, would the common ratio be different? Explain your answer. *I think it would right?

OpenStudy (anonymous):

Because its going to fall faster and bounce higher. So that would change it right?

OpenStudy (anonymous):

If ball 1 were perfect it would have to be the same. But since it isn't, you'll get slight variations...

OpenStudy (anonymous):

any doubts?

OpenStudy (anonymous):

nope! thankyou so much! You really helped a lot!

OpenStudy (anonymous):

Bye.